# Finite-size corrections to the energy spectra of gapless one-dimensional systems in the presence of boundaries

### v1

@@ 4 Local shift of the system size and boundary energy
4.1 Local shift of the system size $L_{\alpha/\beta}$
@@
$L=N$ (110)
as in the PBC case. However, it is noticed that in the presence of the open boundaries, there is $O(1/N^{2})$ FSC in the TF-Ising model, and it can be absorbed into the fourth term in the RHS of (109) formally by simply redefining an effective system size $L_{\text{eff}}\neq N$ as motivated in Ref. [50, 37].
@@
where $L_{\alpha}$ and $L_{\beta}$ are "local shifts of the system size" that are determined only by the specific boundary conditions $\alpha$ and $\beta$ applied on the ends of the spin chain, respectively. Apparently, $\delta L\sim a$. In this way, we claim that to match the lattice and BCFT spectrum, one needs to fix the system size in BCFT to be
$L=(N+(\delta L)_{\alpha}+(\delta L)_{\beta})a\ .$ (114)
@@
$+\sum_{i}a_{n}^{\prime i}(N+(\delta L)_{\alpha}+(\delta L)_{\beta})^{-b_{n}^{\prime i}}\ ,$ (115)
where $\mathfrak{E}_{\alpha/\beta}\coloneqq\varepsilon_{\alpha/\beta}-\varepsilon_{0}(\delta L)_{\alpha/\beta}$ are called the redefined boundary energy and will be discussed further in the next subsection. $\{a^{\prime i}_{n}\}$ and $\{b^{\prime i}_{n}\}$ are the coefficients and powers in expanding $E_{n}^{\alpha\beta}(N)$ as a series of $N+(\delta L)_{\alpha}+(\delta L)_{\beta}$. Here, we remark that, unlike the PBC case, one can only explain these coefficients through the perturbation theory if one chooses the correct $(\delta L)_{\alpha/\beta}$. Otherwise, the set of the coefficients $\{a^{\prime i}_{n}\}$ and $\{b^{\prime i}_{n}\}$ will be a mixture of the Taylor expansion of all preceding perturbations. This feature of the systems with open boundaries will make applying Cardy’s perturbation theory significantly harder than the PBC case.
4.2 Redefinition of boundary energy $E_{\alpha/\beta}$
@@
$\delta\left(\mathcal{E}\right)^{(n)}_{g}\sim\text{const.}\frac{1}{L}\left(\frac{g}{L^{\Delta-2}}\right)^{n}+\text{const.}\frac{1}{a}\left(\frac{g}{a^{\Delta-2}}\right)^{n}+\text{const.}\frac{L}{a^{2}}\left(\frac{g}{a^{\Delta-2}}\right)^{n},$ (117)
@@
Similarly, for a boundary perturbation $g^{L}=g^{R}=g^{B}$, the $n$-th order corrections take the form
$\delta\left(\mathcal{E}\right)^{(n)}_{g^{B}}\sim\text{const.}\frac{1}{L}\left(\frac{g^{B}}{L^{h-1}}\right)^{n}+\text{const.}\frac{1}{a}\left(\frac{g^{B}}{a^{h-1}}\right)^{n},$ (118)
@@ Acknowledgements
Y. L. is supported by the Global Science Graduate Course (GSGC) program of the University of Tokyo. A. U. is supported by the MERIT-WINGS Program at the University of Tokyo, the JSPS fellowship (DC1). He was supported in part by MEXT/JSPS KAKENHI Grants No. JP21J2052. The work of M. O. is supported by JSPS KAKENHI Grant Nos. JP23H01094, JP23K25791 and JP24H00946. This work was also supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP), during the visit there by M. O. and Y. L. A part of the computation in this paper has been done using the facilities of the Supercomputer Center, the Institute for Solid State Physics, the University of Tokyo.

### v2

@@ 4 Local shift of the system size and boundary energy
4.1 Local shift of the system size $\delta L$
@@
$L=Na$ (112)
as in the PBC case. However, it is noticed that in the presence of the open boundaries, there is $O(1/N^{2})$ FSC in the TF-Ising model, and it can be absorbed into the fourth term in the RHS of (111) formally by simply redefining an effective system size $L_{\text{eff}}\neq Na$ as motivated in Ref. [54, 41].
@@
where $L_{\alpha}$ and $L_{\beta}$ are "local shifts of the system size" that are determined only by the specific boundary conditions $\alpha$ and $\beta$ applied on the ends of the spin chain, respectively. Apparently, $\delta L\sim O(1)$. In this way, we claim that to match the lattice and BCFT spectrum, one needs to fix the system size in BCFT to be
$L=\left[N+(\delta L)_{\alpha}+(\delta L)_{\beta}\right]a\ .$ (116)
@@
$+\sum_{i}a_{n}^{\prime i}\left[N+(\delta L)_{\alpha}+(\delta L)_{\beta}\right]^{-b_{n}^{\prime i}}\ ,$ (117)
where $\mathfrak{E}_{\alpha/\beta}\coloneqq\varepsilon_{\alpha/\beta}-\varepsilon_{0}(\delta L)_{\alpha/\beta}$ are called the redefined boundary energy and will be discussed further in the next subsection. $\{a^{\prime i}_{n}\}$ and $\{b^{\prime i}_{n}\}$ are the coefficients and powers in expanding $E_{n}^{\alpha\beta}(N)$ as a series of $N+(\delta L)_{\alpha}+(\delta L)_{\beta}$. Here, we remark that, unlike the PBC case, one can only explain these coefficients through the perturbation theory if one chooses the correct $(\delta L)_{\alpha/\beta}$. Otherwise, the set of the coefficients $\{a^{\prime i}_{n}\}$ and $\{b^{\prime i}_{n}\}$ will be a mixture of the Taylor expansion of all preceding perturbations. This feature of the systems with open boundaries will make applying Cardy’s perturbation theory significantly harder than the PBC case. This ambiguity is non-universal and greatly depends on the concrete realization of the boundary conditions in the lattice model, and there is no general theoretical description of the shift $\delta L$. Nevertheless, as we have seen, sometimes the shifts take simple values and/or obey a certain relation when an exact solution or a Kramers-Wannier type duality is applicable to the lattice model.
4.2 Redefinition of boundary energy $\mathfrak{E}$
@@
$\delta\left(\mathcal{E}\right)^{(n)}_{g}\sim\text{const.}\frac{1}{L}\left(\frac{g}{L^{\Delta-2}}\right)^{n}+\text{const.}\frac{1}{b}\left(\frac{g}{b^{\Delta-2}}\right)^{n}+\text{const.}\frac{L}{b^{2}}\left(\frac{g}{b^{\Delta-2}}\right)^{n},$ (119)
@@
Similarly, for a boundary perturbation $g^{L/R}=g^{B}$, the $n$-th order corrections take the form
$\delta\left(\mathcal{E}\right)^{(n)}_{g^{B}}\sim\text{const.}\frac{1}{L}\left(\frac{g^{B}}{L^{h-1}}\right)^{n}+\text{const.}\frac{1}{b}\left(\frac{g^{B}}{b^{h-1}}\right)^{n},$ (120)
@@ Acknowledgements
Y. L. is supported by the Global Science Graduate Course (GSGC) program of the University of Tokyo. A. U. is supported by the MERIT-WINGS Program at the University of Tokyo, the JSPS fellowship (DC1). He was supported in part by MEXT/JSPS KAKENHI Grants No. JP21J2052. The work of M. O. is supported by JSPS KAKENHI Grant Nos. JP23H01094, JP23K25791 and JP24H00946. This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP), during the visit there by M. O. and Y. L. A part of the computation in this paper has been done using the facilities of the Supercomputer Center, the Institute for Solid State Physics, the University of Tokyo.